Partner Bonus Burns has a capital balance of $118,000 after adjusting assets to fair market value. Murray contributes $64,000 to receive a 35% in a new partnership with Burns Determine the amount and recipient of the partner bonus Amount of bonus Recipient of bonus
media%2Fffa%2Fffa5e136-68b4-4254-b541-a2
media%2F2c9%2F2c9195ef-4bfb-4f6e-94e7-4f
Answer #1
PARTNER BONUS
Old profit sharing ratio of Burns before admission = 100% =1
Total capital after admission = $118000+$64000 =$182000
New profit sharing ratio of burns = $118000 / $ 182000 = 0.65
New profit sharing ratio of Murray = $64000 / $ 182000 = 0.35
Sacrificing ratio of burns = Old ratio - new ratio = 1- 0.65 = 0.35
Amount of bonus = $118000 x 0.35 = $ 41300
Recipient of bonus = Burns
LIDQUATING PARTNERSHIPS
Amount realised on sale of assets = $48000
Amount of liabilites = $1000
Therefore, Net assets = $48000 -$1000 = $47000
Total capital = $16000 +$27000 = $43000
Excess of net assets over capital that is profit will be shared equally among partners = $47000-$43000= $4000
Amount of final distribution on liquation of partnership to MacPherson = share of capital +profits share of net assets
                          = $16000 + half of profits on net assets $4000 = $18000
LIQUIDATING PARTNERSHIPS -DEFICIENCY
Amount realised on sale of assets = $40000
Amount of liabilites = NIL
Therefore, Net assets = $40000 -NIL = $40000
Total capital = $21000 +$79000 = $100000
Excess of capital over net assets that loss / deficiency on liqudation = $ 100000-$40000= $60000
The loss is shared equally = $60000/2 = $30000
Amount of Pepper's deficiney = $21000 -$30000 =($ 9000)
Amount to be distributed to Morrison = $79000 -$30000 = $69000
Amount to be distributed to Morrison assuming Pepper is unable to satisfy the deficiney (Pepper's deficiency is adjusted) = $69000 - $9000 = $60000